# Collaborative Coded Caching for Partially Connected Networks

### v2

@@ IV Numerical Results
We use numerical results to compare the performance of the least cost branch and bound and its greedy approximation and the scalable heuristic of [14] that applies to the collision interference model. As a quick explanation, the heuristic in [14] categorizes the users according to the number of helpers within their transmission radius and assigns them to helpers in a greedy way by avoiding collisions. For the simulation setup, we assume $H$ helpers are located at the center of hexagons on a limited hexagonal grid. Every hexagon has a radius of $1$ (normalized length unit). The users are placed in a circular area with a radius of $r_{u}=2.7$ centered on the center of the hexagonal lattice according to a homogeneous Poisson Point Process, with density per unit area $u$ (see Figure 6). Users are randomly assigned to a cache profile $l\in[L]$. The error bars in the figures correspond to the standard deviation.
In Figure 6, we plot the achievable sum-DoF as a function of $L$. We can observe that the least cost branch and bound method and its greedy approximation outperform the heuristic of [14] as expected since it assumes the model is more restricted, i.e., the collision interference model. Since $u/L$ is fixed, on average, as $L$ changes, the number of users with the same cache profile doesn’t change so for the proposed methods, the number of partitions per cache profile remains constant. And since $\gamma L$ increases linearly as $L$ increases, the number of partitions that can be served simultaneously increases linearly, and as a consequence, we observe a linear increase in the sum-DoF. In Figure 6, we plot the achievable sum-DoF as a function of $r$ together with the optimum sum-DoF under uncoded cache placement for fully connected networks with $C_{\ell}>=R$ for all $\ell\in[L]$ [12]. We can observe that as $r$ increases, the performance of the proposed methods increases; on the other hand, the performance of the heuristics of [14] decreases. The reasoning is as follows: The proposed methods use the spatial multiplexing gain of the transmitters to serve multiple users by finding nonzero links between the transmitters and users. So, as $r$ increases, the network’s connectivity increases, and sum-DoF increases. However, the heuristic of [14] serves the users by avoiding collisions and there are fewer possibilities for collision-free transmissions. Notice that the achieved sum-DoF of the proposed methods are the same when the network becomes fully connected at $r=4.2$ as expected. We also see from Figure 6 that the proposed methods achieve a sum-DoF close to optimal when the network becomes fully connected.

### v3

@@ IV Numerical Results
We provide simulations to compare the performance of the least cost branch and bound and its greedy approximation (see Example 1) and the scalable heuristic of [14] that applies to the collision interference model. As a quick explanation, the heuristic in [14] categorizes users based on the number of helpers within their transmission radius and assigns them to helpers in a greedy way by avoiding collisions. For the simulation setup, we assume $E$ helpers are located at the center of hexagons on a limited hexagonal grid. Every hexagon has a radius of $1$ (normalized length unit). The users are placed in a circular area with a radius of $r_{u}=2.7$ centered on the center of the hexagonal lattice according to a homogeneous Poisson Point Process, with density per unit area $u$ (see Figure 6). Users are randomly assigned to a cache profile $l\in[L]$. The error bars in the figures correspond to the standard deviation.
In Figure 6, we plot the achievable sum-DoF as a function of $L$. We can observe that the least cost branch and bound method and its greedy approximation outperform the heuristic of [14] as expected, since it assumes a more restricted model, i.e., the collision interference model. Since $u/L$ is fixed, on average, as $L$ changes, the number of users with the same cache profile doesn’t change, so, for the proposed methods, the number of partitions per cache profile remains constant. And since $\gamma L$ increases linearly as $L$ increases, the number of partitions that can be served simultaneously increases linearly, and as a consequence, we observe a linear increase in the sum-DoF. In Figure 6, we plot the achievable sum-DoF as a function of $r$ together with the optimum sum-DoF under uncoded cache placement for fully connected networks with $C_{\ell}>=E$ for all $\ell\in[L]$ [12]. We can observe that as $r$ increases, the performance of the proposed methods improves, while the performance of the heuristics of [14] decreases. The reasoning is as follows: The proposed methods use the spatial multiplexing gain of the transmitters to serve multiple users by finding nonzero links between the transmitters and users. As $r$ increases, the network’s connectivity increases, and sum-DoF increases. However, the heuristic of [14] serves the users by avoiding collisions, and there are fewer possibilities for collision-free transmissions. Notice that the achieved sum-DoF of the proposed methods are the same when the network becomes fully connected at $r=4.2$ as expected. We also see from Figure 6 that the proposed methods achieve a sum-DoF close to optimal when the network becomes fully connected.